# ProtoGuard-guided PROPEL: Class-Aware Prototype Enhancement and Progressive Labeling for Incremental 3D Point Cloud Segmentation

### v1

@@ III-B ProtoGuard: Geometric and Semantic Prototype Enhancement
First, the geometric and semantic prototypes are initialized from their respective features. As shown in Fig. 2, the geometric feature $\mathcal{F}_{g}$ is derived from the spatial coordinates ($r,s,t$) of the original point cloud, combined with the normal vectors and height information extracted from EdgeConv3 [32]. Meanwhile, the semantic feature $\mathcal{F}_{s}$ is computed from the density distribution, derived from the spatial coordinates of the original point cloud, along with the original semantic information extracted by EdgeConv3.
Second, we dynamically maintain the prototypes of each class by updating the prototypes since the initial prototype is unstable, which was directly initialized via averaging the individual features of a particular class. To update the prototype, we leverage an adaptive momentum $\alpha$, which controls the smoothness of prototype updates, preventing them from updating too quickly or too slowly. We compute $\alpha$ based on geometric and semantic similarities, which are defined as the cosine similarity between the current class prototype and the features of other classes, formulated as:
$S_{\text{geo }}=\frac{\mathcal{P}_{g}^{\text{cur }}\cdot\mathcal{F}_{g}^{\text{other }}}{\left\|\mathcal{P}_{g}^{\text{cur }}\right\|\left\|\mathcal{F}_{g}^{\text{other }}\right\|},$ (1)
$S_{\text{sem }}=\frac{\mathcal{P}_{s}^{\text{cur }}\cdot\mathcal{F}_{s}^{\text{other }}}{\left\|\mathcal{P}_{s}^{\text{cur }}\right\|\left\|\mathcal{F}_{s}^{\text{other }}\right\|},$
where $\mathcal{P}_{g}^{\text{cur }}$ and $\mathcal{P}_{s}^{\text{cur }}$ are current geometric and semantic prototypes, $\mathcal{F}_{g}^{\text{other }}$ and $\mathcal{F}_{s}^{\text{other }}$ represent features from other classes. With the computed similarities, we obtain adaptive momentum as:
$\alpha_{\mathrm{geo}}=\alpha_{\mathrm{geo}}\cdot\left(1-S_{\mathrm{geo}}\right)+\lambda_{\mathrm{geo}}\cdot S_{\mathrm{geo}},$ (2)
$\alpha_{\mathrm{sem}}=\alpha_{\mathrm{sem}}\cdot\left(1-S_{\mathrm{sem}}\right)+\lambda_{\mathrm{sem}}\cdot S_{\mathrm{sem}},$
where $\alpha_{\mathrm{geo}}$ and $\alpha_{\mathrm{sem}}$ are adaptive momentums, $\lambda_{\text{geo }}$ and $\lambda_{\text{sem }}$ are learnable adjustment factors. The prototypes are then updated using these momentums:
$\mathcal{P}_{g}^{\text{new }}=\alpha_{\text{geo }}\cdot\mathcal{P}_{g}^{\text{cur }}+\left(1-\alpha_{\text{geo }}\right)\cdot\mathcal{F}_{g}^{\text{cur }},$ (3)
$\mathcal{P}_{s}^{\text{new }}=\alpha_{\text{sem }}\cdot\mathcal{P}_{s}^{\text{cur }}+\left(1-\alpha_{\text{sem }}\right)\cdot\mathcal{F}_{s}^{\text{cur }}.$
The geometric and semantic features are finally fused through an attention-based mechanism:
$\mathcal{F}_{\text{enhanced }}$ $=w_{g}\mathcal{F}_{g}+w_{s}\mathcal{F}_{s},$ (4)
with $w_{g}$ and $w_{s}$ being learnable parameters.
Last, our final feature fusion combines prototype features and edge features, followed by MLP and max pooling. The edge features $\mathcal{F}_{\text{edge }}$ are obtained by calculating the feature differences between each pair of neighboring points through the EdgeConv3 layer. To sum up, the final feature can be formulated as follows:
$\mathcal{F}_{\text{final}}=\text{MLP}(\text{MaxPooling}(\mathcal{F}_{\text{enhanced}}\oplus\mathcal{F}_{\text{edge}})),$ (5)
where $\oplus$ denotes channel-wise concatenation. The base classifier ${C}_{b}$ processes these final features to predict point-wise segmentation labels:
$\hat{y}_{i}={C}_{b}(\mathcal{F}_{\text{final }})_{i}.$ (6)
The network is optimized using the cross-entropy segmentation loss:
$L_{\text{base }}=-\sum_{i}\mathcal{L}_{\mathrm{seg}}\left(y_{i},\hat{y}_{i}\right),$ (7)
where $y_{i}$ and $\hat{y}_{i}$ denote the ground truth and predicted labels for point $d_{i}$, respectively.
@@ III-C PROPEL: Progressive Refinement Of PsEudo-Labels
Second, estimate the point-wise uncertainty in novel point clouds. Here, we apply a spatial sampling approach based on Bayesian Active Learning by Disagreement (BALD) [33]. BALD effectively captures distribution uncertainty through multiple forward passes with neighborhood configurations. For each point $d_{n}^{i}$ in the novel point cloud, uncertainty is computed as:
$\mathcal{U}_{n}^{i}=H(E_{t}(p(c|d_{n}^{i},\mathcal{N}_{t})))-E_{t}(H(p(c|d_{n}^{i},\mathcal{N}_{t}))),$ (8)
where $H(\cdot)$ represents information entropy, $p(c|d_{n}^{i},\mathcal{N}_{t})$ denotes the class prediction probability distribution for point $d_{n}^{i}$ under the $t$-th neighborhood configuration $\mathcal{N}_{t}$, and ${E}_{t}(\cdot)$ signifies expectation over multiple sampling instances (i.e., averaging across different neighborhood configurations). It measures the difference between the entropy of the expected prediction and the expected entropy of predictions, quantifying the model’s prediction consistency across different spatial contexts. Specifically, we leverage weighted predictions based on spatial proximity, where weights $\sigma_{n}^{t}$ derive from normalized cosine similarities between target points and their neighbors:
$\mathcal{U}_{n}^{i}=H\left(\frac{1}{T}\sum_{t}Q_{base}^{i,c}\cdot\sigma_{n}^{t}\right)-\frac{1}{T}\sum_{c,t}H(Q_{base}^{i,c}\cdot\sigma_{n}^{t}),$ (9)
where $Q_{base}^{i,c}$ represents the probability distribution predicted by the base classifier for point $d_{n}^{i}$ belonging to class $c$, $T$ is the total number of spatial neighborhood configurations sampled, and $\sigma_{n}^{t}$ weights the predictions according to spatial proximity. This implementation explicitly models the expectation terms in the theoretical equation by averaging over $T$ different neighborhood configurations while accounting for the spatial structure inherent in point cloud data.
@@
Furthermore, we use adaptive thresholding to determine pseudo-label reliability. To address the varying uncertainty levels across different class regions and mitigate errors in pseudo-label propagation, particularly in areas where classes overlap and are prone to misclassification, we develop a context-aware adaptive threshold mechanism that accounts for local point distribution characteristics:
$\tau=\tau_{base}\cdot f_{den}(d_{n}^{i})\cdot f_{sem}(d_{n}^{i})\cdot f_{uncertain}(\mathcal{U}_{n}^{i}),$ (10)
where $\tau_{base}$ is the base threshold value which is set to 0.9, $f_{den}(d_{n}^{i})=(1+\gamma\cdot\rho_{norm}(d_{n}^{i}))$ captures local density variation, with $\rho_{norm}(d_{n}^{i})$ representing the normalized local density and $\gamma$ controlling its influence $f_{sem}(d_{n}^{i})=(1+\lambda\cdot S_{sem}(d_{n}^{i}))$ incorporates semantic relevance to novel classes, where $S_{sem}(d_{n}^{i})$ measures semantic similarity to known novel class features and $\lambda$ is its adjustment weight $f_{uncertain}(\mathcal{U}_{n}^{i})=(1+\beta\cdot\mathcal{U}_{n}^{i})$ accounts for prediction stability, adaptively increasing the threshold in regions with higher uncertainty, with $\beta$ controlling this adjustment.
Our pseudo-label generation follows a hierarchical decision process that selectively integrates knowledge from base and novel models:
[…]
where $\mathbf{H}_{n}^{i}$ represents the generated pseudo-label for point $d_{n}^{i}$, $\mathbf{L}_{n}^{i}$ is the initial label prediction, $y_{bg}$ represents the “background” class outside the base class set $Y_{base}$, and $y_{bg}^{\prime}$ denotes the “background” class beyond the novel class set $Y_{novel}$. The decision process prioritizes reliable predictions from the base classifier while handling uncertain regions through neighborhood information.
This hierarchical approach addresses four key scenarios by: (a) trusting high-confidence base classifier predictions when they disagree with initial background classifications; (b) leveraging reliable neighborhood information when the base classifier is uncertain; (c) preserving existing non-background novel class labels; and (d) excluding ambiguous points to maintain training quality. For neighborhood-based label retrieval, we select neighbor points $q$ that satisfy both non-background prediction ($Q_{base}^{q,y}\neq y_{bg}$) and low uncertainty ($\mathcal{U}_{n}^{q}\leq\tau$). This selective approach deliberately excludes regions where both target points and their neighbors exhibit high uncertainty, as these typically contain noise that would degrade model performance.
The generated pseudo-labels are used to train the novel classifier with the segmentation loss:
$\mathcal{L}_{seg}=-\sum_{i}\sum_{y}\mathbf{H}_{n}^{i}\log(Q_{novel}^{i,\hat{y}}),$ (12)
where $\mathbf{H}_{n}^{i}$ is the pseudo-label vector for point $d_{i}$ and $Q_{novel}^{i,\hat{y}}$ represents the predicted probability distribution from the novel classifier.
@@ IV-A Experimental Settings
Dataset. In our experiments, we utilize two widely used publicly available datasets: S3DIS [12] and ScanNet [11], chosen for their diversity and relevance to the task of point cloud semantic segmentation. S3DIS consists of point clouds from 272 rooms across six different indoor areas. Each point in the dataset includes both XYZ coordinates and RGB color information. The points are annotated with one of 13 predefined semantic classes. Following [27], we select Area 5 as the validation set for our experiments, while the remaining areas are used for training. ScanNet is an RGB-D video dataset that includes 1,513 scans taken from 707 indoor scenes. Each point in the dataset is labeled with one of 21 classes, which include 20 semantic classes and an additional class for unannotated areas. We use the 1,210 scans in the dataset for training, while the remaining 312 scans are reserved for validation. For both datasets, we follow the standard training and validation splits and the number of novel classes as in [27], ensuring a fair comparison with benchmark methods. As shown in Tab I and Tab II, $S^{0}$ represents the class indices consistent with the original dataset, while $S^{1}$ represents the class indices arranged in alphabetical order. $C_{novel}$ is set to 5, 3, and 1, respectively.
Implementation details. In the base-class training phase, we first use the three-layer EdgeConv layers of DGCNN [32] to extract initial features, which our ProtoGuard then refines to generate more reliable features. Together, these form our feature extractor. In the novel-class training phase, our base class feature extractor and classifier are inherited from the previous phase and frozen. The new class weights in the novel-class feature extractor and classifier are randomly initialized. In contrast, the remaining parts are initialized with the weights from the pre-trained model on the base classes. Regarding parameter settings, we follow the configurations in [27]. We train the model with a batch size of 32 and the Adam optimizer [36], with an initial learning rate of 0.001 and a weight decay of 0.0001 for 100 epochs. After 50 epochs, the learning rate is decayed by a factor of 0.5. We also choose the nearest neighbors $K=12$ and $T=12$, and we set the uncertainty threshold $\tau$ to 0.0065 for S3DIS and 0.0045 for ScanNet through cross-validation. The value of $r$ is set to 0.25. All experiments are executed on 4$\times$×NVIDIA GeForce RTX4090.

### v2

@@ III-B ProtoGuard: Geometric and Semantic Prototype Enhancement
First, the geometric and semantic prototypes are initialized from their respective features. As shown in Fig. 2, the geometric feature is combined with the normal vectors and height information. Meanwhile, the semantic feature is computed from the density distribution, derived from the spatial coordinates of the original point cloud, along with the original semantic information.
Second, we dynamically maintain the prototypes of each class by updating the prototypes since the initial prototype is unstable, which was directly initialized via averaging the individual features of a particular class. To update the prototype, we leverage an adaptive momentum, preventing them from updating too quickly or too slowly. We compute based on geometric and semantic similarities, which are defined as the cosine similarity between the current class prototype and the features of other classes. With the computed similarities, we obtain adaptive momentum, which are learnable adjustment factors. The prototypes are then updated using these momentums. The geometric and semantic features are finally fused through an attention-based mechanism.
Last, our final feature fusion combines prototype features and edge features, followed by MLP and max pooling. The edge features are obtained by calculating the feature differences between each pair of neighboring points. The network is optimized using the cross-entropy segmentation loss.
@@ III-C PROPEL: Progressive Refinement Of PsEudo-Labels
Second, estimate the point-wise uncertainty in novel point clouds. Here, we apply a spatial sampling approach based on Bayesian Active Learning by Disagreement (BALD) [32]. BALD effectively captures distribution uncertainty through multiple forward passes with neighborhood configurations. For each point $d_{n}^{i}$ in the novel point cloud, uncertainty is computed. It measures the difference between the entropy of the expected prediction and the expected entropy of predictions, quantifying the model’s prediction consistency across different spatial contexts. Specifically, we leverage weighted predictions based on spatial proximity. This implementation explicitly models the expectation terms in the theoretical equation by averaging over $T$ different neighborhood configurations while accounting for the spatial structure inherent in point cloud data.
@@
Furthermore, we use adaptive thresholding to determine pseudo-label reliability. To address the varying uncertainty levels across different class regions and mitigate errors in pseudo-label propagation, particularly in areas where classes overlap and are prone to misclassification, we develop a context-aware adaptive threshold mechanism that accounts for local point distribution characteristics. Our pseudo-label generation follows a hierarchical decision process that selectively integrates knowledge from base and novel models. The decision process prioritizes reliable predictions from the base classifier while handling uncertain regions through neighborhood information.
This hierarchical approach addresses four key scenarios by: (a) trusting high-confidence base classifier predictions when they disagree with initial background classifications; (b) leveraging reliable neighborhood information when the base classifier is uncertain; (c) preserving existing non-background novel class labels; and (d) excluding ambiguous points to maintain training quality. For neighborhood-based label retrieval, we select neighbor points $q$ that satisfy both non-background prediction ($Q_{base}^{q,y}\neq y_{bg}$) and low uncertainty ($\mathcal{U}_{n}^{q}\leq\tau$). This selective approach deliberately excludes regions where both target points and their neighbors exhibit high uncertainty, as these typically contain noise that would degrade model performance. The generated pseudo-labels are used to train the novel classifier with the segmentation loss.
@@ IV-A Experimental Settings
Dataset. In our experiments, we utilize two widely used publicly available datasets: S3DIS [12] and ScanNet [11], chosen for their diversity and relevance to the task of point cloud semantic segmentation. S3DIS consists of point clouds from 272 rooms across six different indoor areas. Each point in the dataset includes both XYZ coordinates and RGB color information. The points are annotated with one of 13 predefined semantic classes. Following [27], we select Area 5 as the validation set for our experiments, while the remaining areas are used for training. ScanNet is an RGB-D video dataset that includes 1,513 scans taken from 707 indoor scenes. Each point in the dataset is labeled with one of 21 classes, which include 20 semantic classes and an additional class for unannotated areas. We use the 1,210 scans in the dataset for training, while the remaining 312 scans are reserved for validation. For both datasets, we follow the standard training and validation splits and the number of novel classes as in [27], ensuring a fair comparison with benchmark methods.